OpenStudy (anonymous):

Use the associative law of addition to write an equivalent expression. (7 + u) + 5v

Parth (parthkohli):

Associative law looks something like: \(\Large \color{Black}{\Rightarrow (a + b) + c = a + (b + c) \Longleftrightarrow (7 + u) + 5v = 7+(u + 5v)}\)

OpenStudy (anonymous):

Thank you